[diffusion] Diffusion Morning Report 06/27/13

Ted Berg tberg at stanford.edu
Thu Jun 27 06:12:59 PDT 2013


Good Morning All,
     Epi Chamber B pump out for repair.

Diffusion Morning ReportJune 27, 2013

*New Problems: :AMAT Epi B Pressure , *

*Ongoing: SEM 4160 tilt issue, ,*

*Resolved: Epi B scrubber pump*

*Tools*

**

	

*Status*

**

	

*Priority*

	

*Problem/Comment*

	

*Action Plan/Update*

*AMAT Epi Cham B*

	

Down

	

1

	

Pressure issues ,

	

Send out pump

*SEM 4160*

	

Yellow

	

2

	

Tilt issues

	

Check out-maybe FS

*Tylan 11-LTO*

	

Yellow

	

	

LEDs

	

*Thermco1*

	

Yellow

	

	

Thin

	

*Thermco BPSG*

	

Yellow

	

	

Front Zone low

	

Check element


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20130627/f2a550d6/attachment-0001.html>


More information about the diffusion mailing list